OpenStudy (anonymous):

Solve the following system of equations. 2x + y = 3 x = 2y - 1 @hartnn

OpenStudy (anonymous):

Just plug in the second expression into the x value of the first expression so you will get \[2(2y-1) + y = 3\] then solve for y, plug in that value into either equation and solve for x.

jimthompson5910 (jim_thompson5910):

Hint 2x + y = 3 2(2y-1) + y = 3 ... plug in x = 2y - 1 4y - 2 + y = 3 5y - 2 = 3 I'll let you finish up

OpenStudy (anonymous):

y=1

jimthompson5910 (jim_thompson5910):

now use that to find x

OpenStudy (anonymous):

x=1

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 2x + y = 6 y = 3x + 4 What is the resulting equation? y+ 2x+y= 6 2x+y+ 3x+ 4 = 6 2x+ (3x+ 4) = 6

jimthompson5910 (jim_thompson5910):

which one do you think it is? use the last problem as a guide

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

y=16

jimthompson5910 (jim_thompson5910):

but what is the answer to the initial question

OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 3x + y = 1 y + 4 = 5x What is the resulting equation?

jimthompson5910 (jim_thompson5910):

you need to solve for y in the second equation first

OpenStudy (anonymous):

Based on the lesson, which of the following would be the best approach for solving this system by substitution? 5x = y + 6 2x - 3y = 4 Solve the first equation forx. Solve the first equation fory. Solve the second equation forx. Solve the second equation fory.

jimthompson5910 (jim_thompson5910):

which one do you think it is